OpenStudy (anonymous):

How would solve 3n^4-4n^2+1=0 by factoring?

OpenStudy (anonymous):

substitute t = n^2

OpenStudy (anonymous):

My teacher wants us to use either factoring or the quadratic formula on these problems.

OpenStudy (anonymous):

Yes. first you have to change it to quadratic then you can factor

OpenStudy (anonymous):

okay let me see...

OpenStudy (anonymous):

kayyy

OpenStudy (anonymous):

use this hint \[\large 3n^4-4n^2+1=3n^4-3n^2-(n^2-1)=3n^2(n-1)-(n-1)(n+1)\]

OpenStudy (anonymous):

Yes grouping gives quick factors here. check the steps of @mukushla

OpenStudy (anonymous):

thank y'all.